# INTERVENOR[Uncaptioned image]: Prompting the Coding Ability of Large Language Models with the Interactive Chain of Repair

### v4

@@ Abstract
This paper introduces INTERVENOR (INTERactiVE chaiN Of Repair), a system designed to emulate the interactive code repair processes observed in humans, encompassing both code diagnosis and code repair. INTERVENOR prompts Large Language Models (LLMs) to play distinct roles during the code repair process, functioning as both a Code Learner and a Code Teacher. Specifically, the Code Learner is tasked with adhering to instructions to generate or repair code, while the Code Teacher is responsible for crafting a Chain-of-Repair (CoR) to serve as guidance for the Code Learner. During generating the CoR, the Code Learner needs to check the generated codes from Code Learner and reassess how to address code bugs based on error feedback received from compilers. Experimental results demonstrate that INTERVENOR surpasses baseline models, exhibiting improvements of approximately 18% and 4.3% over GPT-3.5 in code generation and code translation tasks, respectively. Our further analyses show that CoR is effective to illuminate the reasons behind bugs and outline solution plans in natural language. With the feedback of code compilers, INTERVENOR can accurately identify syntax errors and assertion errors and provide precise instructions to repair codes. All data and codes are available at https://github.com/NEUIR/INTERVENOR.
@@
[…]
@@ 2 Related Work
Another research avenue involves employing an iterative code repair approach to enhance the quality of generated code Zhang et al. (2023a); Welleck et al. (2023); Madaan et al. (2023); Shinn et al. (2023). Self-Debug Chen et al. (2023b) utilizes explanations generated by LLMs to rectify self-generated code, while Self-Repair Olausson et al. (2023) incorporates human-provided feedback for improvement. Self-Edit Zhang et al. (2023a) employs error messages to refine generated code, but it necessitates the training of an additional fault-aware editor to generate a new program. Instead of directly integrating feedback for code repair, INTERVENOR designs an additional agent to reflect the reasons for coder errors and generate the Chain-of-Repair (CoR). This CoR is then employed to instruct the other agent in code repair through natural language.
@@ 3.2.2 Interactive Code Repair Workflow
[…]
@@ 4 Experimental Methodology
HumanEval Chen et al. (2021) serves as a benchmark for evaluating the functional correctness of synthesized programs generated from docstrings. It comprises 164 hand-written Python programming problems, which consist of function signatures, docstrings, bodies, and multiple unit tests. MBPP Austin et al. (2021) is a benchmark that includes 974 introductory-level Python programming problems. Each problem comprises a problem statement, a code solution, and three automated test cases, and the task IDs that range from 11 to 510 are used for evaluation. HumanEval-X Zheng et al. (2023) is used to assess a model’s multi-programming language generation and translation capability. It consists of 820 human-crafted data instances, covering C++, Java, JavaScript, and Go.
@@
Implementation Details. In our experiments, we use GPT-3.5 as the foundation model to build different agents in our INTERVENOR model. We set the temperature to 0.2 and the maximum generation length to 512 tokens. The maximum number of interactive code repairs is set to 5. Additionally, we also use CodeLlama-7B/13B to implement the agents, Code Learner and Code Teacher, of our INTERVENOR model to explore the impact of using different LLMs.
@@ 5 Evaluation Results
Model Code Generation Code Translation
HumanEval MBPP HumanEval-X Target Language
Python Python C++ Java JS Python C++ Java JS
InCoder Fried et al. (2023) 15.2 19.4 9.5 9.1 13.0 - - - -
CodeGen Nijkamp et al. (2023) 18.3 20.9 18.1 14.9 18.4 40.7 37.6 35.4 51.8
CodeGeeX Zheng et al. (2023) 22.9 24.4 17.1 20.0 17.6 68.5 43.6 56.8 45.2
CodeT5+ Wang et al. (2023b) 30.9 - - - - - - - -
InstructCodeT5+ Wang et al. (2023b) 35.0 - - - - - - - -
PaLM-Coder Chowdhery et al. (2023) 35.9 47.0 - - - - - - -
StarCoder Li et al. (2023b) 40.8 49.5 - - - - - - -
WizardCoder Luo et al. (2023) 57.3 51.8 - - - - - - -
LLama2 Touvron et al. (2023) 30.5 45.4 - - - - - - -
CodeLLama Roziere et al. (2023) 62.2 61.2 - - - - - - -
PanGu-Coder2 Shen et al. (2023) 61.6 - - - - - - - -
Claude Anthropic (2023) 47.6 - - - - - - - -
GPT-4 OpenAI (2023) 67.0 - - - - - - - -
Self-Debug Chen et al. (2023b) 73.8 - - - - - - - -
Self-Collaboration Dong et al. (2023) 74.4 68.2 - - - - - - -
GPT-3.5 OpenAI (2022) 60.3 39.8 52.4 50.6 54.3 84.3 71.5 81.7 84.6
INTERVENOR 75.6 69.8 67.1 68.3 67.1 89.8 75.6 85.4 88.3
Table 2: Overall Performance of Different Models. We evaluate model effectiveness on code generation and code translation (HumanEval-X dataset) tasks using the Pass@$1$ evaluation metric. The baseline results are borrowed from corresponding papers. More evaluation results on the code translation task are shown in Appendix A.2.
@@ 5.1 Overall Performance
Code Repair Prompt Methods HumanEval MBPP HumanEval-X CodeError Avg.
@@
Table 3: Ablation Studies. We compare different prompting techniques to evaluate the effectiveness of our Chain-of-Repair (CoR) mechanism. All models are built based on GPT-3.5 and evaluated with Pass@$1$.
Code Learner Code Teacher HumanEval MBPP HumanExal-X CodeError
@@ 5.2 Ablation Studies
Firstly, we evaluate the code generation ability of LLMs using different code generation/repair prompting methods. As shown in Table 3, we compare CoT Kojima et al. (2022), Few-Shot Chen et al. (2021), and Few-Shot CoT Wei et al. (2022) models in experiments, which prompt LLMs to better generate codes. These models try to generate natural language as the chain of coding thought (CoT) or provide some instances to demonstrate the coding task (Few-Shot). Then we compare different methods to generate the code repair instruction, including Self-Refine Madaan et al. (2023), Error Msgs, and Chain-of-Repair (CoR). Self-Refine asks the LLMs to rethink the errors by themselves, while Error Msgs and CoR incorporate the code error messages from compilers. Msgs directly uses error messages to guide the code repair process.
@@
Then, as shown in Table 4, we investigate the impact of using different LLMs to build Code Learner and Code Teacher. When we use the identical LLM to build the Code Learner, the performance of INTERVENOR is improved with the enhanced capabilities of the Code Teacher, underscoring the efficacy of our CoR mechanism. This indicates that stronger models can conduct a more in-depth analysis of erroneous code and provide more accurate suggestions for code repair. Similarly, by keeping the same LLM to build the Code Teacher, a stronger Code Learner also conducts better code repair results. This indicates that a more powerful Code Learner has a stronger instruction-following ability to better comprehend the CoR provided by the Code Teacher.
@@ 6 Conclusion
This paper proposes INTERVENOR, which regards the feedback of compilers as the INTERVENOR to facilitate the agent collaboration of code error diagnose and repair. Our experiments show that INTERVENOR outperforms previous code generation/repair models by generating accurate code repair solutions and instructions.

### v5

@@ Abstract
This paper introduces INTERVENOR (INTERactiVE chaiN Of Repair), a system designed to emulate the interactive code repair processes observed in humans, encompassing both code diagnosis and code repair. INTERVENOR prompts Large Language Models (LLMs) to play distinct roles during the code repair process, functioning as both a Code Learner and a Code Teacher. Specifically, the Code Learner is tasked with adhering to instructions to generate or repair code, while the Code Teacher is responsible for crafting a Chain-of-Repair (CoR) to serve as guidance for the Code Learner. During generating the CoR, the Code Teacher needs to check the generated codes from Code Learner and reassess how to address code bugs based on error feedback received from compilers. Experimental results demonstrate that INTERVENOR surpasses baseline models, exhibiting improvements of approximately 18% and 4.3% over GPT-3.5 in code generation and code translation tasks, respectively. Our further analyses show that CoR is effective to illuminate the reasons behind bugs and outline solution plans in natural language. With the feedback of code compilers, INTERVENOR can accurately identify syntax errors and assertion errors and provide precise instructions to repair codes. All data and codes are available at https://github.com/NEUIR/INTERVENOR.
@@
Hanbin Wang1, Zhenghao Liu1thanks:   indicates corresponding author., Shuo Wang2, Ganqu Cui2, Ning Ding2, Zhiyuan Liu2 and Ge Yu1 1Department of Computer Science and Technology, Northeastern University, China 2Department of Computer Science and Technology, Institute for AI, Tsinghua University, China Beijing National Research Center for Information Science and Technology, China
@@ 2 Related Work
Another research avenue involves employing an iterative code repair approach to enhance the quality of generated code Zhang et al. (2023a); Welleck et al. (2023); Madaan et al. (2023); Shinn et al. (2023); Josifoski et al. (2024). Self-Debug Chen et al. (2023b) utilizes explanations generated by LLMs to rectify self-generated code, while Self-Repair Olausson et al. (2023) incorporates human-provided feedback for improvement. Self-Edit Zhang et al. (2023a) employs error messages to refine generated code, but it necessitates the training of an additional fault-aware editor to generate a new program. Instead of directly integrating feedback for code repair, INTERVENOR designs an additional agent to reflect the reasons for coder errors and generate the Chain-of-Repair (CoR). This CoR is then employed to instruct the other agent in code repair through natural language.
@@ 3.2.2 Interactive Code Repair Workflow
In the initial step (Step A0, where 00 signifies the initial turn), we prompt the Code Learner to generate code for the given task. Subsequently, the Code Learner executes the generated code using the code compiler to assess its correctness (Step B). Following this, the Code Teacher generates code repair instructions (CoR) based on the bug report and the associated buggy code (Step C). These instructions elucidate the reason of the bug, such as “modified to handle the case where the input decimal is 0”, and include code correction planning, for example, “we can simply return db0db”. Such instructions are informative and enhance the guidance for the Code Learner. Ultimately, the Code Learner follows the chain-of-repair (CoR) to rectify the code and subsequently resubmits the corrected version to the compiler for execution in the subsequent turn ((Step Ai), where $i\geq 1$ denotes the code repair process). The Ai, B, and C steps are iterated sequentially until either the code meets the compiler’s estimation or the maximum turn limit is reached.
@@ 4 Experimental Methodology
HumanEval (HEval) Chen et al. (2021) serves as a benchmark for evaluating the functional correctness of synthesized programs generated from docstrings. It comprises 164 hand-written Python programming problems, which consist of function signatures, docstrings, bodies, and multiple unit tests. MBPP Austin et al. (2021) is a benchmark that includes 974 introductory-level Python programming problems. Each problem comprises a problem statement, a code solution, and three automated test cases, and the task IDs that range from 11 to 510 are used for evaluation. HumanEval-X Zheng et al. (2023) is used to assess a model’s multi-programming language generation and translation capability. It consists of 820 human-crafted data instances, covering C++, Java, JavaScript, and Go. DS-1000 Lai et al. (2022) is a benchmark designed to evaluate the capabilities of LLMs in data science code generation. It includes 1,000 problems that span seven Python libraries, such as NumPy and Pandas. We use the completion style prompt for each question.
@@
Implementation Details. In our experiments, we use GPT-3.5 (gpt-3.5-turbo-0613) as the foundation model to build different agents in INTERVENOR. We set the temperature to 0.2 and the maximum generation length to 512 tokens. The maximum number of interactive code repairs is set to 5. Additionally, we also use CodeLlama-7B/13B-Instruct to implement the agents, Code Learner and Code Teacher, of our INTERVENOR model to explore the impact of using different LLMs. On all datasets, we use the 0-shot setting in our experiments.
Model Code Generation Code Translation
HEval MBPP HumanEval-X DS-1000 Target Language
Python Python C++ Java JS Python Python C++ Java JS
InCoder Fried et al. (2023) 15.2 19.4 9.5 9.1 13.0 7.4 - - - -
CodeGen Nijkamp et al. (2023) 18.3 20.9 18.1 14.9 18.4 8.4 40.7 37.6 35.4 51.8
CodeGeeX Zheng et al. (2023) 22.9 24.4 17.1 20.0 17.6 - 68.5 43.6 56.8 45.2
CodeT5+ Wang et al. (2023b) 30.9 - - - - - - - - -
InstructCodeT5+ Wang et al. (2023b) 35.0 - - - - - - - - -
PaLM-Coder Chowdhery et al. (2023) 35.9 47.0 - - - - - - - -
StarCoder Li et al. (2023b) 40.8 49.5 - - - 26.0 - - - -
WizardCoder Luo et al. (2023) 57.3 51.8 - - - 28.4 - - - -
LLama2 Touvron et al. (2023) 30.5 45.4 - - - - - - - -
CodeLLama Roziere et al. (2023) 62.2 61.2 - - - 28.0 - - - -
PanGu-Coder2 Shen et al. (2023) 61.6 - - - - - - - - -
Claude Anthropic (2023) 47.6 - - - - - - - - -
GPT-4 OpenAI (2023) 67.0 - - - - - - - - -
Self-Debug (Simple) Chen et al. (2023b) 73.8 - - - - - - - - -
Self-Debug (UT+Trace) Chen et al. (2023b) 71.9 - - - - - - - - -
Self-Collaboration Dong et al. (2023) 74.4 68.2 - - - - - - - -
GPT-3.5 OpenAI (2022) 60.3 39.8 52.4 50.6 54.3 29.7 84.3 71.5 81.7 84.6
INTERVENOR 75.6 69.8 67.1 68.3 67.1 39.7 89.8 75.6 85.4 88.3
Table 2: Overall Performance of Different Models. We evaluate model effectiveness on code generation and code translation (HumanEval-X dataset) tasks using the Pass@$1$ evaluation metric. The baseline results are borrowed from corresponding papers. Simple and UT+Trace are two variants of Self-Debug. More evaluation results on the code translation task are shown in Appendix A.2.
@@ 5.1 Overall Performance
Code Repair Prompt Methods HEval MBPP HumanEval-X CodeError Avg.
@@
Table 3: Evaluations on Different Prompting Methods. We compare INTERVENOR with different prompting techniques to evaluate its effectiveness. All models are built based on GPT-3.5 and evaluated using Pass@$1$.
Code Learner Code Teacher HEval MBPP HumanExal-X CodeError
@@ 5.2 Ablation Studies
Evaluation on Different Prompting Methods. Firstly, we evaluate the code generation ability of LLMs using different code generation/repair prompting methods. As shown in Table 3, we compare CoT Kojima et al. (2022), Few-Shot Chen et al. (2021), and Few-Shot CoT Wei et al. (2022) models in experiments, which prompt LLMs to better generate codes. These models try to generate natural language as the chain of coding thought (CoT) or provide some instances to demonstrate the coding task (Few-Shot). Then we compare different methods to generate the code repair instruction, including Self-Refine Madaan et al. (2023), Error Msgs, and Chain-of-Repair (CoR). Self-Refine asks the LLMs to rethink the errors by themselves, while Error Msgs and CoR incorporate the code error messages from compilers. Msgs directly uses error messages to guide the code repair process.
@@
Building Code Teacher/Learner Using Different LLMs. Then, as shown in Table 4, we investigate the impact of using different LLMs to build Code Learner and Code Teacher. When we use the identical LLM to build the Code Learner, the performance of INTERVENOR is improved with the enhanced capabilities of the Code Teacher, underscoring the efficacy of our CoR mechanism. This indicates that stronger models can conduct a more in-depth analysis of erroneous code and provide more accurate suggestions for code repair. Similarly, by keeping the same LLM to build the Code Teacher, a stronger Code Learner also conducts better code repair results. This indicates that a more powerful Code Learner has a stronger instruction-following ability to better comprehend the CoR provided by the Code Teacher.
#Turn Model HEval MBPP HumanEval-X Avg.
Python Python C++ Java JS
0 GPT-3.5 60.3 39.8 52.4 50.6 54.3 51.5
1 Error Msgs 67.1 51.8 57.3 59.8 62.8 59.8
Self-Refine 65.2 48.8 57.3 64.0 60.4 59.1
INTERVENOR 69.5 51.0 59.8 65.2 60.4 61.2
2 Error Msgs 68.3 53.0 60.3 62.8 63.4 61.6
Self-Refine 69.5 49.6 60.3 64.6 62.8 61.4
INTERVENOR 73.2 54.4 61.6 67.1 62.2 63.7
3 Error Msgs 69.5 54.9 62.2 63.7 64.0 62.9
Self-Refine 71.9 51.2 63.4 65.2 63.4 63.0
INTERVENOR 75.6 60.2 65.2 68.3 64.0 66.7
Table 5: Multi-Turn Code Repair Performance of INTERVENOR, Error Msgs, and Self-Refine.
Effectiveness on Alleviating the DoT problem. Finally, we conduct experiments to demonstrate how our model alleviates the Degenerate-of-Thought (DOT) problem. As shown in Table 5, we compare the multi-turn code repair performance of INTERVENOR, Error Msgs, and Self-Refine. Both INTERVENOR and Error Msgs incorporate feedback from the compiler, while Self-Refine conducts self-reflection to generate the reasons for buggy code. Thus, Self-Refine usually faces the DoT problem, which usually neglects the bugs that appear in the code segments. The experiments show that with the increase in the number of repair turns, the performance of INTERVENOR surpasses that of Self-Refine and Error Msgs across almost all datasets. After three turns of code repair, INTERVENOR achieves an average improvement of 3.7% over Self-Refine and Error Msgs, highlighting its effectiveness in alleviating the DoT problem.
@@ 6 Conclusion
This paper proposes INTERVENOR, which regards the feedback of compilers as the INTERVENOR to facilitate the agent collaboration of code error diagnosis and repair. Our experiments show that INTERVENOR outperforms previous code generation/repair models by generating accurate code repair solutions and instructions.
@@
Acknowledgments
Hanbin Wang and Zhenghao Liu are supported by the Natural Science Foundation of China under Grant (No. 62206042 and No. 62272093), the Joint Funds of Natural Science Foundation of Liaoning Province (No. 2023-MSBA-081), and the Fundamental Research Funds for the Central Universities under Grant (No. N2416012). Ge Yu is supported by the Natural Science Foundation of China under Grant (No. 62137001).
@@ A.3 More Details of the CodeError Benchmark
Turns Self-Debug Self-Collaboration Self-Refine INTERVENOR (Ours)
1 3 >3 2 2
2 6 >6 4 4
3 9 >9 6 6
Table 10: A Comparison of Computational Overhead Between INTERVENOR and Other Iterative Self-Refinement Baselines.
@@ A.5 Prompts for Agent Building
A.6 Computational Overhead
In this subsection, we compare the computational overhead with other iterative self-refinement baselines.
We follow Self-Debug, Self-Collaboration, and Self-Refine to iteratively fix buggy codes. As shown in Table 10, if we iterate for the same number of turns, INTERVENOR calls the API fewer times and brings performance improvements.